# Meta-Transfer Derm-Diagnosis: Exploring Few-Shot Learning and Transfer Learning for Skin Disease Classification in Long-Tail Distribution

### v1

@@ 4 The Methodology
This section outlines the core methodologies deployed in our study, starting with the Meta-Transfer Derm-Diagnosis Framework. This framework is pivotal for assessing four distinct model training strategies that are particularly valuable in scenarios with limited data samples. These strategies are Few-Shot Episodic Transfer Learning (FETL), Few-Shot Episodic Learning (FEL), Deep Transfer Learning (DTL), and Standard Deep Learning (DL).
@@ 4.1 Meta-Transfer Derm-Diagnosis Framework
Definition 1. For the training and testing phases, the novel dataset $D_{\text{novel}}$ is split into two subsets: the support set ($D_{S}$) and the query set ($D_{Q}$). In a typical few-shot classification scenario, the support set $D_{S}$ contains only a few samples per class, often ranging from 1 to 5. The primary goal in few-shot classification is to train a classifier, $f:X_{\text{novel}}\rightarrow Y_{\text{novel}}$, using the limited data in $D_{S}$. This classifier should then be able to accurately categorize instances in the query set $D_{Q}$. If $D_{S}$ includes $N$ distinct classes with $K$ labeled examples per class, this scenario is defined as an $N$-way $K$-shot classification. The case with only one labeled example per class is termed one-shot classification.
@@
To further clarify our methodology, we define the datasets used. The base dataset is defined as $D_{\text{base}}=\{(x_{i},y_{i});x_{i}\in X_{\text{base}},y_{i}\in Y_{\text{base}}\}_{i=1}^{N_{\text{base}}}$, where $x_{i}$ represents the feature vector of the $i$-th image, and $y_{i}$ is its corresponding class label. The novel dataset is similarly represented as $D_{\text{novel}}=\{(\tilde{x}_{j},\tilde{y}_{j});\tilde{x}_{j}\in X_{\text{novel}},\tilde{y}_{j}\in Y_{\text{novel}}\}_{j=1}^{N_{\text{novel}}}$. It is crucial to note that the class labels in $D_{\text{base}}$ and $D_{\text{novel}}$ are mutually exclusive, i.e., $Y_{\text{base}}\cap Y_{\text{novel}}=\emptyset$. In a similar manner, $D_{\text{imgnet}}$ is represented as: $\{(\bar{x}_{k},\bar{y}_{k});\bar{x}_{k}\in X_{\text{imgnet}},\bar{y}_{k}\in Y_{\text{imgnet}}\}_{k=1}^{N_{\text{imgnet}}}$.
To rigorously evaluate the framework we devised, detailed in Figure 3, we maintained a consistent evaluation by keeping $D_{\text{novel}}$ fixed. Here, $D_{\text{train}}$ denotes the datasets used during training. We formulated four distinct training methodologies, each independently designed: FETL, where $D_{\text{train}}=D_{\text{imagenet}}+D_{\text{base}}*$ ; FEL, where $D_{\text{train}}=D_{\text{base}}*$; DTL, where $D_{\text{train}}=D_{\text{imagenet}}+D_{\text{base}}$; and DL, where $D_{\text{train}}=D_{\text{imagenet}}$. The notation $D_{\text{base}}*$ indicates episodic training, while the others involve traditional large-scale training approaches for the corresponding datasets in that domain.
In accordance with these definitions, as is summarized in Figure 4, our FETL, FEL, and DTL models utilize $D_{\text{base}}$ and $D_{\text{novel}}$ classes from the same domain for training and testing as specified in Definition 1. On the other hand, the proposed DL model utilizes two separate domains for training and testing, hence it aligns with Definition 2. Therefore, three out of our four proposed analysis include adapted few-shot training and testing methodologies considering the data extracted carefully from the same long-tail distributions (FETL, FEL and DTL), while the last one (DL) corresponds to a cross-domain evaluation.
Algorithm 1 $N$-Way $K$-Shot Classification Evaluation.
0:  $D_{\text{train}}=\{(x_{i},y_{i});X_{i}\in\mathcal{X}_{\text{train}},Y_{i}\in\mathcal{Y}_{\text{train}}\}_{i=1}^{N_{\text{train}}}.$
0:  $D_{\text{novel}}=\{(\tilde{x}_{j},\tilde{y}_{j});\tilde{x}_{j}\in\mathcal{X}_{\text{novel}},\tilde{y}_{j}\in\mathcal{Y}_{\text{novel}}\}_{j=1}^{N_{\text{novel}}}.$
@@
4:     Randomly select $M$ samples from the remaining samples of $N$ classes as the query set $\{(\tilde{x}^{(e)},\tilde{y}^{(e)})\}$.
@@ 4.3 Regularization via Image Augmentation
Regularization techniques are essential in preventing overfitting and enhancing the generalization capabilities of deep models. Among these techniques, image augmentation stands out as a crucial method in supervised learning, well-known for its efficacy in regularization. While conventional augmentation methods like rotation, horizontal flips, and vertical flips are widely used, they often prove inadequate in domains characterized by limited data, such as medical imaging. Consequently, advanced approaches such as MixUp [50], CutMix [51], and ResizeMix [52] have received significant attention due to their ability to address the challenges posed by data scarcity and variability.These methods provide advanced solutions that extend beyond conventional techniques, facilitating the generation of diverse training samples and enhancing model robustness from various perspectives.
MixUp
MixUp generates synthetic samples by blending features and labels from two different images using weighted blending. Let $(x_{i},y_{i})$ and $(x_{j},y_{j})$ represent the features and labels of two images, respectively. MixUp combines these two images with weighted blending to create a new synthetic sample:
$\tilde{x}=\lambda x_{i}+(1-\lambda)x_{j}$
$\tilde{y}=\lambda y_{i}+(1-\lambda)y_{j}$
Here, $\lambda\in[0,1]$ is a random weight value.
CutMix
CutMix creates synthetic samples by replacing regions of an image with patches from another image, controlled by a cut ratio ($\lambda$). Let $(x_{i},y_{i})$ and $(x_{j},y_{j})$ denote the features and labels of two images, respectively. CutMix blends these two images according to the cut ratio to create a new synthetic sample:
$\tilde{x}=M\cdot x_{i}+(1-M)\cdot x_{j}$
$\tilde{y}=\lambda y_{i}+(1-\lambda)y_{j}$
Here, $M$ is a mask matrix, and $\lambda$ is a random cut ratio.
ResizeMix
ResizeMix merges images of different sizes by resizing them and then blending them using weighted blending. Let $(x_{i},y_{i})$ and $(x_{j},y_{j})$ represent the features and labels of two images, respectively. ResizeMix resizes one image to the size of the other and blends them with weighted blending to create a new synthetic sample:
$\tilde{x}=\lambda x_{i}+(1-\lambda)\text{resize}(x_{j},\text{size}(x_{i}))$
$\tilde{y}=\lambda y_{i}+(1-\lambda)y_{j}$
Here, $\text{resize}(x_{j},\text{size}(x_{i}))$ resizes $x_{j}$ to the size of $x_{i}$, and $\lambda\in[0,1]$ is a random weight value.
@@ 5 Experiments and Discussion
Figure 5: The graph compares the accuracies of the models based on the number of shots. It illustrates the results for a) SD-198, b) Derm7pt, and c) ISIC2018.
Table 1: Performance comparison of the proposed models on the SD-198 skin lesion dataset for 2-way and 5-way classification tasks. Accuracy (%) values are presented.
Exp MobileNetV2 DenseNet121
FEL FETL DTL DL FEL FETL DTL DL
2W-1S 82.66 84.77 82.42 77.77 82.91 82.83 81.21 75.70
2W-2S 87.41 89.57 88.34 85.98 87.77 88.26 87.88 84.39
2W-3S 89.77 91.73 91.43 89.52 90.23 90.82 91.38 88.63
2W-4S 90.72 93.05 93.38 91.44 91.90 92.47 93.16 90.48
2W-5S 91.49 93.66 94.40 92.96 92.95 93.58 94.56 91.86
2W-6S 91.85 93.81 95.09 93.65 93.63 94.19 95.26 92.58
2W-7S 92.61 94.41 95.73 94.68 94.20 94.75 96.02 93.71
2W-8S 92.85 94.79 96.21 95.35 94.55 95.22 96.67 94.27
2W-9S 93.02 94.89 96.46 95.48 94.79 95.56 96.88 94.64
2W-10S 93.19 95.10 96.77 95.69 95.05 95.74 97.07 94.94
5W-1S 61.73 65.85 64.06 56.93 63.61 64.81 62.61 54.25
5W-2S 69.93 75.02 75.72 71.07 72.15 73.95 75.30 68.56
5W-3S 73.97 79.04 80.88 76.81 76.97 78.47 80.73 74.57
5W-4S 76.47 81.48 84.32 80.88 80.06 81.95 84.53 78.51
5W-5S 78.16 83.22 86.67 83.24 82.05 84.02 86.95 81.03
5W-6S 78.84 83.97 88.10 85.31 83.62 85.59 88.78 83.16
5W-7S 79.76 85.04 89.17 86.52 84.68 86.60 90.22 84.58
5W-8S 80.36 85.50 90.09 87.47 85.52 87.35 91.05 85.55
5W-9S 81.14 85.94 90.83 88.55 86.30 88.32 91.91 86.50
5W-10S 81.40 86.36 91.37 89.35 86.77 88.51 92.57 87.33
Table 2: Performance comparison of the proposed models on the Derm7pt skin lesion dataset for 2-way and 5-way classification tasks. Accuracy (%) values are presented.
Exp MobileNetV2 DenseNet121
FEL FETL DTL DL FEL FETL DTL DL
2W-1S 59.37 61.38 59.65 58.65 61.40 60.99 60.00 59.42
2W-2S 63.87 66.20 65.15 64.53 65.83 66.13 67.45 65.46
2W-3S 66.45 69.22 68.67 68.60 69.21 70.21 71.17 69.88
2W-4S 68.21 71.07 72.27 71.36 70.83 72.24 74.31 72.99
2W-5S 69.18 72.34 75.41 74.15 72.46 74.12 77.25 75.89
2W-6S 70.37 73.57 77.07 75.97 73.34 75.21 79.23 77.56
2W-7S 71.11 74.69 78.81 77.29 74.50 76.83 80.90 79.15
2W-8S 71.40 74.82 80.07 78.44 74.66 76.67 81.69 79.98
2W-9S 72.25 75.81 81.59 79.87 75.21 77.98 83.54 80.99
2W-10S 71.98 76.38 82.43 80.34 75.68 78.28 83.90 81.68
5W-1S 32.38 32.79 31.80 31.70 31.78 33.74 32.04 31.80
5W-2S 39.16 40.39 41.06 40.65 39.73 41.54 42.78 42.34
5W-3S 41.97 43.96 46.89 46.51 43.35 46.44 49.19 48.67
5W-4S 43.88 45.87 51.82 51.19 45.64 49.92 53.77 53.57
5W-5S 45.38 47.81 55.92 54.69 47.05 52.48 57.68 56.83
5W-6S 46.02 48.83 58.67 57.47 47.87 53.96 60.92 59.56
5W-7S 47.24 50.31 61.42 59.97 48.86 55.65 63.65 61.45
5W-8S 47.70 50.71 63.63 61.78 49.31 56.87 65.93 63.22
5W-9S 48.16 51.49 65.12 63.20 49.81 57.46 67.80 64.51
5W-10S 48.58 52.15 66.33 64.38 50.11 58.56 69.28 65.48
Table 3: Performance comparison of the proposed models on the ISIC2018 skin lesion dataset for 2-way and 3-way classification tasks. Accuracy (%) values are presented.
Exp MobileNetV2 DenseNet121
FEL FETL DTL DL FEL FETL DTL DL
2W-1S 58.42 58.49 64.83 60.55 57.86 59.04 58.55 60.66
2W-2S 62.62 62.23 73.10 67.99 61.55 62.92 66.60 68.46
2W-3S 63.99 63.86 77.26 71.69 63.63 64.98 71.16 72.34
2W-4S 65.38 65.49 79.81 74.02 64.83 66.55 74.31 74.77
2W-5S 66.94 66.31 81.63 76.42 65.97 67.51 76.69 76.40
2W-6S 67.65 67.01 82.55 77.07 66.17 68.27 78.31 77.24
2W-7S 68.05 67.73 83.52 78.22 67.04 68.49 79.30 78.08
2W-8S 68.08 67.73 84.02 78.55 67.68 68.81 80.27 78.68
2W-9S 68.83 68.13 84.75 79.57 67.54 69.35 80.72 79.45
2W-10S 69.50 68.82 85.18 80.39 68.58 70.07 81.99 80.14
3W-1S 42.33 43.09 49.51 45.61 42.08 42.25 43.63 44.63
3W-2S 45.64 45.85 59.10 52.83 45.44 45.89 52.61 52.89
3W-3S 48.02 48.33 63.98 57.75 47.35 48.68 57.70 57.67
3W-4S 49.65 49.39 67.13 60.55 48.80 50.05 61.93 60.65
3W-5S 50.41 50.92 69.35 63.22 50.39 51.56 64.47 62.72
3W-6S 51.43 51.46 71.06 65.06 50.81 52.49 66.50 64.20
3W-7S 52.04 52.06 72.47 66.54 51.28 53.02 68.27 65.44
3W-8S 52.97 52.56 73.16 67.33 52.09 53.53 69.32 65.99
3W-9S 52.91 52.90 74.21 68.31 52.41 53.61 70.60 67.21
3W-10S 53.76 53.21 75.01 69.07 53.07 53.46 71.38 67.59
@@ 5.1 Implementation Details
The implementation of our study is carried out using the Python programming language with the PyTorch library. Nvidia 1080Ti GPU is employed during the model development. To ensure a fair comparison with the Meta-DermDiagnosis [8] model, we aligned the distinctions between base and novel classes in the datasets and ensured the similarity of deleted classes. All datasets are resized to 224x224x3 dimensions. ISIC2018 consists of 4 base and 3 novel classes, Derm7pt includes 13 base and 5 novel classes, and SD-198 encompasses 128 base and 70 novel classes. In contrast to the Meta-DermDiagnosis study, we partitioned the SD-198 base classes into training and validation sets. Classes with fewer than 20 data points are designated as novel, while those with 20-30 data points are assigned to the validation set. To ensure a standardized testing environment for the four differently trained models, we employed a seed and deterministic mode, which worked in allowing each model to be tested on tasks of the same difficulty and order. Hyperparameters such as batch size used during testing were structured similarly for FETL and FEL models, with the only difference being the query set size set to 5 due to data insufficiency in novel classes. Data augmentation techniques are not applied to novel classes. All experiments are conducted using MobileNetV2 and DenseNet121 backboned models.
The sole distinction between the FEL (Few-Shot Episodic Learning) and FETL (Few-Shot Episodic Transfer Learning) models lies in the use of pretrained ImageNet weights for the backbone models in the FETL model. In contrast, the FEL model is trained on backbone models with random initialization. Data augmentation techniques such as RandomResizedCrop, RandomFlip, and ColorJitter are employed. In this framework, our models are trained with a 5-way 5-shot configuration, and all layers of the backbone models are fully opened for training.
Both DTL (Deep Transfer Learning) and DL (Deep Learning) models utilize ImageNet weights, but the DTL model is fine-tuned with the base classes of the datasets. The DL model, serving as a baseline, is chosen to compare performance without any fine-tuning, using ImageNet weights. In the DTL model, traditional training methods are employed instead of episodic training. For instance, in the case of the SD-198 dataset, 10% of the 128-class base section are allocated as the validation set. The MobileNetV2 model with ImageNet weights are then trained, aiming for the model to learn generalized features from the dataset. During testing, this model is used with ProtoNet in an episodic structure to evaluate and compare its performance.
@@
5.2 Experimental Analysis
In order to evaluate our framework and assess the effectiveness of our proposed methods, we performed experiments using 2-Way 1-Shot to 10-Shot and 5-Way 1-Shot to 10-Shot setups. These tests are conducted using three datasets: SD-198, Derm7pt, and ISIC2018. The comprehensive results of our model tests are detailed in Tables 1, 2, and 3. Moreover, we executed additional experiments based on the parameters specified in Tables 4, 5, and 6. These experiments are designed to compare our model’s performance against benchmarks set in prior research, such as SCAN [19], MetaMed [18], and PFEMed [25], using the datasets referred to in Section [9, 10, 11]. In all these tables, the highest accuracy values are highlighted in bold, and the second highest results are underlined.
Table 4: Comparison of our models and the SOTA methods. Values in the table are F1-scores of the corresponding models on the SD-198 dataset.

### v2

@@ 4 The Methodology
This section outlines the core methodologies deployed in our study, starting with the Meta-Transfer Derm-Diagnosis Framework. This framework is pivotal for assessing five distinct model training strategies that are particularly valuable in scenarios with limited data samples. These strategies include Few-Shot Episodic Transfer Learning (FETL), Few-Shot Episodic Learning (FEL), Deep Transfer Learning (DTL), Deep Transfer Learning with Self-Supervised Adaptation (DTLS), and Standard Deep Learning (DL).
@@ 4.1 Meta-Transfer Derm-Diagnosis Framework
Definition 1. For the training and testing phases, the novel dataset $D_{\text{novel}}$ is split into two subsets: the support set ($D_{S}$) and the query set ($D_{Q}$). In a typical few-shot classification scenario, the support set $D_{S}$ contains only a few samples per class, often ranging from 1 to 5. The primary goal in few-shot classification is to train a classifier, $f:X_{\text{novel}}\rightarrow Y_{\text{novel}}$, using the limited data in $D_{S}$. The classifier should then be able to accurately categorize the instances in the query set $D_{Q}$. If $D_{S}$ includes $N$ distinct classes with $K$ labeled examples per class, this scenario is defined as an $N$-way $K$-shot classification. The case with only one labeled example per class is termed one-shot classification.
@@
To further clarify our methodology, we define the datasets used. The base dataset is defined as
$D_{\text{base}}=\{(x_{i},y_{i});x_{i}\in X_{\text{base}},y_{i}\in Y_{\text{base}}\}_{i=1}^{N_{\text{base}}}$, where $x_{i}$ represents the feature vector of the $i$-th image, and $y_{i}$ is its corresponding class label. The novel dataset is similarly represented as
$D_{\text{novel}}=\{(\tilde{x}_{j},\tilde{y}_{j});\tilde{x}_{j}\in X_{\text{novel}},\tilde{y}_{j}\in Y_{\text{novel}}\}_{j=1}^{N_{\text{novel}}}$. It is crucial to note that the class labels in $D_{\text{base}}$ and $D_{\text{novel}}$ are mutually exclusive, i.e., $Y_{\text{base}}\cap Y_{\text{novel}}=\emptyset$. In a similar manner, $D_{\text{imgnet}}$ is represented as: $\{(\bar{x}_{k},\bar{y}_{k});\bar{x}_{k}\in X_{\text{imgnet}},\bar{y}_{k}\in Y_{\text{imgnet}}\}_{k=1}^{N_{\text{imgnet}}}$.
To rigorously evaluate the framework we devised, detailed in Figure 4, we maintained a consistent evaluation by keeping $D_{\text{novel}}$ fixed. Here, $D_{\text{train}}$ denotes the datasets used during training. We formulated five distinct training methodologies, each independently designed: FETL, where $D_{\text{train}}=D_{\text{imagenet}}+D_{\text{base}}*$ ; FEL, where $D_{\text{train}}=D_{\text{base}}*$; DTL and DTLS, where $D_{\text{train}}=D_{\text{imagenet}}+D_{\text{base}}$; and DL, where $D_{\text{train}}=D_{\text{imagenet}}$. The notation $D_{\text{base}}*$ indicates episodic training, while the others involve traditional large-scale training approaches for the corresponding datasets in that domain.
In accordance with these definitions, as summarized in Figure 5, our FETL, FEL, DTLS and DTL models utilize $D_{\text{base}}$ and $D_{\text{novel}}$ classes from the same domain for training and testing as specified in Definition 1. On the other hand, the proposed DL model utilizes two separate domains for training and testing; hence it aligns with Definition 2. Therefore, three out of our five proposed analyses include adapted few-shot training and testing methodologies considering the data extracted carefully from the same long-tail distributions (FETL, FEL, DTL and DTLS), while the last one (DL) corresponds to a cross-domain evaluation.
Algorithm 1 NN-Way $K$-Shot Classification Evaluation.
[…]
@@
4:  Randomly select $M$ samples from the remaining samples of NN classes as the query set $\{(\tilde{x}^{(e)},\tilde{y}^{(e)})\}$.
@@ 4.3 Regularization via Image Augmentation
Regularization is vital for preventing overfitting and improving the generalization of deep models. Among these, image augmentation is particularly effective in supervised learning. However, common techniques such as rotation or flipping are often insufficient in data-limited domains like medical imaging. To address this, advanced methods such as MixUp [74], CutMix [75], and ResizeMix [76] have been proposed, enabling the generation of more diverse samples and improving model robustness.
@@ 5 Experiments and Discussion
Table 1: Few-shot classification accuracies of backbone models without fine-tuning on SD-198, Derm7pt, and ISIC2018 datasets.
Methods
Params
Flops
SD-198 Derm7pt ISIC2018
2W1S 2W5S 5W5S 2W1S 2W5S 5W5S 2W1S 2W5S 3W5S
MedCLIP(ViT) [49] 27.94 4.5 59.28 64.42 36.6 57.88 67.71 31.16 50.74 51.76 35.09
MedCLIP(ResNet50) [49] 25.56 4.11 66.58 80.56 60.06 57.66 64.62 38.25 51.96 55.64 39.5
BioVIL(ResNet-ViT) [48] 25.13 4.2 61.51 72.25 47.96 56.46 62.52 36.8 50.95 52.82 36.49
ConvNeXt(Base) [40] 88.59 15.36 80.32 92.86 81.15 60.25 76.84 55.95 59.55 72.27 58.57
EfficientNetB5 [41] 30.39 10.8 72.88 85.56 70.12 56.0 64.45 40.05 62.14 69.91 54.35
ResNet50 [36] 25.56 4.12 76.77 92.39 82.15 58.45 74.21 55.06 59.97 76.49 63.61
EfficientNetB4 [41] 19.34 4.66 74.59 88.44 74.33 56.55 68.78 44.58 62.91 72.56 57.0
DenseNet121 [37] 7.98 2.88 75.7 91.86 81.03 59.42 75.89 56.83 60.66 76.4 62.72
MobileNetV2 [38] 3.5 0.32 77.77 92.96 83.24 58.65 74.15 54.69 60.55 76.42 63.22
MoCo-v3(ResNet50) [43] 68.01 4.11 69.06 89.52 78.08 58.82 76.53 58.43 58.6 76.58 64.52
SimSiam(ResNet50) [45] 38.2 4.11 71.28 90.35 79.11 58.81 76.14 58.57 60.55 78.05 65.74
SwAV(ResNet50) [46] 28.35 4.11 70.24 89.57 77.8 57.69 74.95 56.64 58.51 76.88 64.88
SimCLR(ResNet50) [42] 27.97 4.11 71.55 89.9 77.93 58.26 74.47 56.34 59.14 75.11 62.14
MoCo-v3(ViT) [43] 215.68 17.58 75.58 85.77 70.08 88.4 92.18 65.24 53.13 55.98 40.18
DINOv2(ViT) [44] 86.58 152.0 70.44 79.03 57.63 86.89 90.07 57.97 54.02 58.04 42.21
ViT-MMF [47] 86.57 17.58 71.69 87.65 73.74 81.92 91.05 75.62 53.37 58.93 43.46
ViT(Base) [39] 86.57 17.58 81.34 92.55 82.14 86.85 93.05 77.9 52.87 58.28 42.31
Table 2: Few-shot classification accuracies of fine-tuned models on SD-198, Derm7pt, and ISIC2018 datasets.
Methods SD-198 Derm7pt ISIC2018
2W1S 2W5S 2W10S 5W1S 5W5S 2W1S 2W5S 2W10S 5W1S 5W5S 2W1S 2W5S 2W10S 3W1S 3W5S
ResNet50+Aug DTL 78.53 94.47 97.52 59.52 87.24 61.62 78.86 86.71 34.03 60.41 61.7 78.28 83.47 46.72 65.86
ResNet50(SimCLR)+Aug DTLS 82.56 93.00 94.97 62.01 79.20 63.91 80.58 85.62 36.52 63.26 62.9 78.54 82.88 47.58 66.1
ResNet50(SimSiam)+Aug DTLS 71.28 90.35 95.08 47.09 76.01 58.81 76.14 83.13 31.19 58.57 60.65 78.05 82.40 45.05 65.74
ResNet50(MoCo-v3)+Aug DTLS 69.09 89.52 94.88 43.92 76.00 58.82 76.53 83.36 31.15 58.43 58.6 76.58 81.04 43.32 64.52
DenseNet121 FEL 82.91 92.95 95.69 63.61 82.05 61.4 72.46 75.68 31.78 47.05 57.86 65.97 68.58 42.08 50.39
DenseNet121 FETL 82.83 93.58 95.05 64.81 84.02 60.99 74.12 78.28 33.74 52.48 59.04 67.51 70.07 42.25 51.56
DenseNet121 DTL 81.21 94.56 95.74 62.61 86.95 60.0 77.25 83.9 32.04 57.68 58.55 76.69 81.99 43.63 64.47
DenseNet121+Aug DTL 82.29 95.46 97.97 65.54 88.76 62.56 81.00 87.16 34.75 61.91 56.67 65.06 67.85 39.81 48.66
MobileNetV2 FEL 82.66 91.49 93.19 61.73 78.16 59.37 69.18 71.98 32.38 45.38 58.42 66.94 69.5 42.33 50.41
MobileNetV2 FETL 84.77 93.66 95.10 65.85 83.22 61.38 72.34 76.38 32.79 47.81 58.49 66.31 68.82 43.09 50.92
MobileNetV2 DTL 82.42 94.4 96.77 64.06 86.67 59.65 75.41 82.43 31.8 55.92 64.83 81.63 85.18 49.51 69.35
MobileNetV2+Aug DTL 84.63 95.46 97.34 67.94 88.49 60.37 77.41 84.93 33.54 59.45 65.55 83.21 86.83 50.40 71.28
ViT(Base)+LoRA+Aug DTL 82.04 92.80 95.80 64.76 85.60 86.35 93.68 95.56 66.33 80.74 61.59 65.64 67.98 45.61 49.55
@@ 5.1 Implementation Details
Our study was implemented in Python using the PyTorch library on an Nvidia 1080Ti GPU. For fair comparison with Meta-DermDiagnosis [9], we matched the base/novel class splits and resized all datasets to 224×224×3. ISIC2018 includes 4 base and 3 novel classes, Derm7pt has 13 base and 5 novel classes, and SD-198 contains 128 base and 70 novel classes. Unlike Meta-DermDiagnosis, we further split SD-198 base classes: those with fewer than 20 samples were set as novel, and those with 20–30 as validation. For consistency, all models were trained with fixed seeds and deterministic settings, ensuring tasks of equal difficulty and order. Hyperparameters such as batch size used during testing were structured similarly for FETL and FEL models, with the only difference being the query set size set to 5 due to data insufficiency in novel classes. Data augmentation techniques are not applied to novel classes. A diverse set of backbone architectures, including convolutional and transformer-based models, was used throughout the experiments, reflecting both conventional and self-supervised transfer learning settings.
The sole distinction between the FEL and FETL models lies in the use of pretrained ImageNet weights for the backbone models in the FETL model. In contrast, the FEL model is trained on backbone models with random initialization. Data augmentation techniques such as RandomResizedCrop, RandomFlip, and ColorJitter are employed. In this framework, our models are trained with a 5-way 5-shot configuration, and all layers of the backbone models are fully opened for training.
Similarly, both DTL, DTLS and DL models utilize ImageNet weights, but the DTL and DTLS model is fine-tuned with the base classes of the datasets. The DL model, serving as a baseline, is chosen to compare performance without any fine-tuning, using ImageNet weights. In the DTL and DTLS setups, a standard training approach is utilized instead of episodic training; for instance, in SD-198, 10% of the base classes are set aside for validation during fine-tuning. DTLS models were pretrained in a self-supervised manner using contrastive loss, and subsequently evaluated by attaching a linear classification head and fine-tuning the models on base classes. The results of this evaluation are presented in Table 2. For models based on Vision Transformer (ViT-Base), fine-tuning was performed using Low-Rank Adaptation (LoRA) to enable efficient parameter updates. For evaluation consistency, all models,including DTL, DTLS, and DL,were tested using a prototypical episodic structure, mimicking few-shot scenarios during inference.
@@
Table 3: Performance comparison of our models and SOTA methods on the SD-198 dataset, reported in F1-scores.